OpenStudy (anonymous):

What is the slope of the line 4x + 6y = 12?

OpenStudy (camerondoherty):

well the formula is Y=Mx+B so what do you think it is?

OpenStudy (muzzack):

\(\huge\ 6y=-4x+12\) \(\huge\ y=\frac{-2}{3}+2\)

OpenStudy (anonymous):

thank you

OpenStudy (muzzack):

your welc ;)